OpenStudy (moonlitfate):

lim (2-3x)/(x^2-1)^(1/2) x→∞

OpenStudy (anonymous):

break it up into [2/(x^2-1)^(1/2)]+[-3x/(x^2-1)^(1/2)] as x approaches infinity, the first term becomes 0. the second term you can further break into -3*[x/(x^2-1)^(1/2)] as x approaches infinity, the stuff on the right side of the times sign becomes infinity/infinity. This becomes 1. therefore, you are left with -3.

OpenStudy (anonymous):

\[\lim_{x\to\infty}\frac{2-3x}{\sqrt{x^2-1}}\\ \lim_{x\to\infty}\frac{2-3x}{\sqrt{x^2}\sqrt{1-\frac{1}{x^2}}}\\ \lim_{x\to\infty}\frac{2-3x}{|x|\sqrt{1-\frac{1}{x^2}}}\]Since x→∞, you have |x|=x, so \[\lim_{x\to\infty}\frac{2-3x}{x\sqrt{1-\frac{1}{x^2}}}\\ \lim_{x\to\infty}\frac{\frac{2}{x}-3}{\sqrt{1-\frac{1}{x^2}}}\]

OpenStudy (anonymous):

\[\large \begin{split} \lim_{x\to \infty }\frac{ 2-3x}{(x^2-1)^{1/2}} &= \lim_{x\to \infty }\frac{ 2-3x}{(x^2-1)^{1/2}} \frac{x^{-1}}{x^{-1}} \\ &=\lim_{x\to \infty }\frac{ 2x^{-1}-3}{(x^{-2})^{1/2}(x^2-1)^{1/2}}\\ &=\lim_{x\to \infty }\frac{ 2x^{-1}-3}{(1-x^{-2})^{1/2}} \\ &=\frac{0-3}{(1-0)^{1/2}} \end{split} \]

OpenStudy (anonymous):

Basically when dealing with infinite limits, a good method is to divide by the highest degree term, keeping in mind that terms inside a square root have their power halved.